Suppose FCFE = $900,000 for years 1-4 and then is expected to grow at a rate of 3%. Assume ke = 18%. If there are 500,000 shares outstanding, what is the predicted price of this stock?
Answer #1

Price of stock is $ 11.22

As per discounted cash flow method, current share price is the present value of cash flows from share.
Step-1:Present value of cash flows of 4 years
Present value of cash flow = Annual Cash flow * Present value of annuity of 1
= $          9,00,000 * 2.690061805
= $ 24,21,055.62
Working:
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.18)^-4)/0.18 i = 18%
= 2.690061805 n = 4
Step-2:Present value of cash flows after 4 years
Present value = CF4*(1+g)/(Ke-g)*DF4 Where,
= $ 31,87,575.25 CF4 = $ 9,00,000
g = 3%
Ke = 18%
DF4 = 1.18^-4 = 0.5157889
Step-3:Price of stock
Price of stock = Sum of present value of cash flows / Shares outstanding
= $ 56,08,630.87 /              5,00,000
= $                11.22
Working:
Sum of present value of cash flows = $ 24,21,055.62 + $ 31,87,575.25
= $ 56,08,630.87